# A Context for Manifold Calculus

### v2

@@ 3.1. Statement of the Main Result
is an equivalence of $\mathcal{C}$. According to [Lur24, Tag 02MD] and [Lur24, Tag 02N9], there is a small category $\mathcal{I}$ and a final functor $N\left(\mathcal{I}\right)\to\mathcal{I}^{\prime}$ whose pullbacks are all final. Then the composite $\chi\circ p$ satisfies the Weiss $k$-condition, so it suffices to consider the case where $\mathcal{I}^{\prime}=N\left(\mathcal{I}\right)$.
Set $n=\dim M$. Using Lemma 3.13, we may assume that $F$ extends to a functor $\mathcal{M}\mathsf{fld}_{n/M}\to\mathcal{C}$ which is left Kan extended from its restriction to $\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}$. By Proposition 2.15, it suffices to show that the functor
$f:\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}\overrightarrow{\times}_{\mathcal{M}\mathsf{fld}_{n/M}}N\left(\mathcal{I}\right)\to\left(\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}\right)^{/\iota}$
is final, where $\iota:U\hookrightarrow M$ denotes the inclusion. By Remark 2.14, the square
${{N(\operatorname{Disj}^{\leq k}(M))\overrightarrow{\times}_{N(\operatorname{Open}(M))}N(\mathcal{I})}}$${{N(\operatorname{Open}(M))^{/U}}}$${{\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}\overrightarrow{\times}_{\mathcal{M}\mathsf{fld}_{n/M}}N(\mathcal{I})}}$${{(\mathcal{D}\mathsf{isk}_{n/M}^{\leq k})^{/\iota}}}$$\scriptstyle{g}$$\scriptstyle{f}$$\scriptstyle{g^{\prime}}$$\scriptstyle{f^{\prime}}$
commutes up to natural equivalence. Since final maps have the right cancellation property [Cis19, Corollary 4.1.9], it will suffice to show that the functors $g$ and $g^{\prime}f^{\prime}$ are final functors.
We start by showing that that $g$ is final. By Propositions 2.12 and 2.23 (and using the equivalence of the ordinary and alternative slices [Lur09, Proposition 4.2.1.5]), it suffices to show that for each object $I\in\mathcal{I}$, the functor
$N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)_{/U_{I}}\to\left(\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}\right)_{/\iota_{I}}$
is final, where $\iota_{I}:U_{I}\hookrightarrow M$ denotes the inclusion. Since the forgetful functor $\left(\mathcal{D}\mathsf{isk}_{n/M}\right)_{/\iota_{I}}\to\mathcal{D}\mathsf{isk}_{n/U_{I}}$ is a trivial fibration, we are reduced to showing that the functor
$N\left(\operatorname{Disj}^{\leq k}\left(U_{I}\right)\right)\to\mathcal{D}\mathsf{isk}_{n/U_{I}}^{\leq k}$
is final. This is a consequence of Lemma 3.14.
Next, we show that the composite $g^{\prime}f^{\prime}$ is final. Set $\mathcal{X}=N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)\overrightarrow{\times}_{N\left(\operatorname{Open}\left(M\right)\right)}N\left(\mathcal{I}\right)$. By the equivalence of the ordinary and alternative slice, it suffices to show that the functor
$\mathcal{X}\to\mathcal{D}\mathsf{isk}_{n/U}^{\leq k}$
defined by $\left(V,I\right)\mapsto\left(V\hookrightarrow U\right)$ is final. For this, we use Lemma 3.14: It suffices to show that, for each finite subset $S\subset U$ of cardinality $\leq k$, the full subcategory $\mathcal{X}_{S}\subset\mathcal{X}$ spanned by the objects $\left(V,I\right)\in\mathcal{X}$ such that $S\subset V$ is weakly contractible. Consider the functor
$p:\mathcal{X}_{S}\to N\left(\mathcal{I}_{S}\right).$
The functor $p$ is a cocartesian fibration whose fibers are the nerves of cofiltered posets. It follows from [Lur09, Proposition 4.1.2.15] that $p$ is an initial functor. In particular, $p$ is a weak homotopy equivalence. Since $N\left(\mathcal{I}_{S}\right)$ is weakly contractible by hypothesis, this shows that $\mathcal{X}_{S}$ is weakly contractible. This completes the proof of (1$\implies$(2).
We now complete the proof by showing that (4$\implies$(1). Let $G$ be a left Kan extension of $F|N\left(\operatorname{Disj}\left(M\right)\right)$, and let $\alpha:G\to F$ be a natural transformation which extends the identity natural transformation of $F|N\left(\operatorname{Disj}\left(M\right)\right)$. We must show that $\alpha$ is a natural equivalence.
Call an open set $U\subset M$ good if the map $\alpha_{U}$ is an equivalence. We wish to show that every open set of $M$ is good. Since $F$ and $G$ are both $k$-excisive and exhaustive (for we already know that (1$\implies$(4)), the collection of good open subsets of $M$ satisfies the hypotheses of Lemma 3.16. Therefore, every open set of $M$ is good, and we are done. ∎
@@
Clearly (2$\implies$(3). Also, we can prove (1$\implies$(2) in exactly the same way as we proved the implication (1$\implies$(2) of Theorem 3.10. It will therefore suffice to show that (3$\implies$(1).
Suppose that $F$ is a Weiss cosheaf. Let $G$ be a left Kan extension of $F|N\left(\operatorname{Disj}\left(M\right)\right)$. Note that $G$ is a Weiss cosheaf because we have already shown that (1$\implies$(3). Let $\alpha:G\to F$ be a natural transformation which extends the identity natural transformation of $F|N\left(\operatorname{Disj}\left(M\right)\right)$. We must show that $\alpha$ is a natural equivalence.
@@
Let $U\subset M$ be an open set homeomorphic to an open subset of $\mathbb{R}^{n}$, where $n=\dim M$. We claim that $U$ is good. Choose an embedding $U\hookrightarrow\mathbb{R}^{n}$ and identify $U$ with a subset of $\mathbb{R}^{n}$. Let $\operatorname{Cube}\left(U\right)$ denote the poset of open subsets of $U$ that are finite disjoint unions of open cubes lying in $U$, and let $\overline{\operatorname{Cube}\left(U\right)}$ denote the poset of open subsets of elements of $\operatorname{Cube}\left(U\right)$. Since $\operatorname{Cube}\left(U\right)$ is closed under finite intersection, the inclusion $N\left(\operatorname{Cube}\left(U\right)\right)\to N\left(\overline{\operatorname{Cube}\left(U\right)}\right)$ is final. Moreover, $\overline{\operatorname{Cube}\left(U\right)}$ is a covering sieve of $U$ in the Weiss topology on $\operatorname{Open}\left(M\right)$. Since every element of $\operatorname{Cube}\left(U\right)$ is good, it follows that $\alpha_{U}$ is an equivalence.
@@
3.4. Taylor Towers and Their Convergence
Now that we have completed the proof of Theorem 3.4, let us use it to construct the Taylor (co)tower of an isotopy functor and discuss when the tower converges to the original functor.
Let $M$ be a manifold and let $\mathcal{C}$ be an $\infty$-category with small colimits. Theorem 3.4 says that for each isotopy functor $F:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{C}$ and each $k\geq 0$, there is a best approximation of $F$ by $k$-excisive, exhaustive isotopy functor, namely, the left Kan extension $T_{k}F$ of $F|N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$. This is the $k$th polynomial approximation of $F$. We now organize these approximations into a cotower
@@ 3.4. Taylor Towers and Their Convergence
Let $\mathcal{A}$ be an $\infty$-category equipped with a sequence $\mathcal{A}^{0}\subset\mathcal{A}^{1}\subset\cdots$ of full subcategories. Set $\mathcal{A}^{\infty}=\bigcup_{i}\mathcal{A}^{i}$. We let $\mathbb{T}\left(\mathcal{A}\right)\subset N\left(\mathbb{Z}_{\geq 0}\cup\{\infty\}\right)\times\mathcal{A}$ denote the full subcategory spanned by the objects $\left(i,A\right)$, where $i\in\mathbb{Z}_{\geq 0}$ and $A\in\mathcal{A}_{i}$, together with the objects $\left(\infty,A\right)$, where $A\in\mathcal{A}$.
Let $\mathcal{C}$ be another $\infty$-category with small colimits, and let $F:\mathcal{A}\to\mathcal{C}$ be a functor. A functor $\mathbb{T}\left(F\right):N\left(\mathbb{Z}_{\geq 0}\cup\{\infty\}\right)\times\mathcal{A}\to\mathcal{C}$ is called a Taylor cotower of $F$ (with respect to the subcategories $\{\mathcal{A}^{i}\}_{i\geq 0}$) if it is a left Kan extension of the composite
[…]
For each $k\in\mathbb{Z}_{\geq 0}\cup\{\infty\}$, we set $T_{k}F=\mathbb{T}\left(F\right)|\{k\}\times\mathcal{A}$. Note that for each $k\in\mathbb{Z}_{\geq 0}$ and $A\in\mathcal{A}$, the inclusion
@@
A Taylor cotower $\mathbb{T}\left(F\right)$ of $F$ is said to be convergent if the functor $N\left(\mathbb{Z}_{\geq 0}\cup\{\infty\}\right)\to\operatorname{Fun}\left(\mathcal{A},\mathcal{C}\right)$ adjoint to $\mathbb{T}\left(F\right)$ is a colimit diagram.
@@
Let $M$ be a manifold, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{C}$ be a functor. A Taylor cotower of $F$ is a Taylor cotower of $F$ with respect to the subcategories $\{\operatorname{Disj}^{\leq i}\left(M\right)\}_{i\geq 0}$. If $F$ is an isotopy functor, then for each $k\in\mathbb{Z}$, the $k$th level $T_{k}F$ of a Taylor cotower of $F$ is polynomial of degree $\leq k$ (Theorem 3.4).
Warning 3.20.
Despite the name, the constituents of Taylor cotowers of non-isotopy functors may fail to be polynomial. For example, consider the functor $F:N\left(\operatorname{Open}\left(\mathbb{R}\right)\right)\to N\left(\mathsf{Set}\right)$ defined by $F\left(U\right)=U\times U$. It is easy to check that $T_{1}F$ is not excisive.
@@
Let $\mathcal{A}$ be an $\infty$-category and let $\mathcal{A}_{0}\subset\mathcal{A}_{1}\subset\cdots$ be a sequence of full subcategories of $\mathcal{A}$. Set $\mathcal{A}_{\infty}=\bigcup_{i\geq 0}\mathcal{A}_{i}$. Let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:\mathcal{A}\to\mathcal{C}$ be a functor. The following conditions are equivalent:
@@
We will use the transitivity of Kan extensions [Lur09, Proposition 4.3.2.8]. Let $T=\mathbb{T}\left(F\right)$ be a Taylor cotower of $F$. Since colimits in functor categories can be formed pointwise [Lur24, Tag 02X9], an easy finality argument shows that condition (1) is equivalent to the following condition:
@@
    $T$ is a left Kan extension of $T|N\left(\mathbb{Z}_{\geq 0}\right)\times\mathcal{A}$.
Let $\mathcal{X}\subset N\left(\mathbb{Z}_{\geq 0}\cup\{\infty\}\right)\times\mathcal{A}$ denote the full subcategory spanned by the objects in $N\left(\mathbb{Z}_{\geq 0}\right)\times\mathcal{A}$ and the objects in {}×$\{\infty\}\times\mathcal{A}_{\infty}$. Since every object $A\in\mathcal{A}_{\infty}$ belongs to some $\mathcal{A}_{i}$, the functor $T|\mathcal{X}$ is a left Kan extension of $T|N\left(\mathbb{Z}_{\geq 0}\right)\times\mathcal{A}$. Therefore, condition (1) is equivalent to the following condition:
@@
On the other hand, $T|\mathcal{X}$ is a left Kan extension of $T|N\left(\mathbb{Z}_{\geq 0}\cup\{\infty\}\right)\times\mathcal{A}_{\infty}$, so (1′′) is equivalent to the following condition:
@@
    $T$ is a left Kan extension of $T|N\left(\mathbb{Z}_{\geq 0}\cup\{\infty\}\right)\times\mathcal{A}_{\infty}$.
Since $T|N\left(\mathbb{Z}_{\geq 0}\right)\times\mathcal{A}$ is a left Kan extension of $T|N\left(\mathbb{Z}_{\geq 0}\right)\times\mathcal{A}_{\infty}$, condition (1′′′) is equivalent to condition (2), and we are done. ∎
@@
Let $M$ be a manifold and let $\mathcal{C}$ be an $\infty$-category with small colimits. An isotopy functor $F:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{C}$ has a convergent Taylor cotower if and only if $F$ satisfies the equivalent conditions of Theorem 3.11.

### v3

@@ 1. Polynomial Functors
is an equivalence of $\mathcal{C}$. According to [Lur25, Tag 02MD] and [Lur25, Tag 02N9], there is a small category $\mathcal{I}$ and a final functor $\mathcal{I}\to\mathcal{I}^{\prime}$ whose pullbacks are all final. Then the composite $\chi\circ p$ satisfies the Weiss $k$-condition, so it suffices to consider the case where $\mathcal{I}^{\prime}=\mathcal{I}$.
Set $n=\dim M$. Using Theorem A.1 (and the fact that localizations are final), we can identify $\theta$ with the map
$\operatorname{colim}_{\left(I,V\right)\in\int^{I\in\mathcal{I}}\mathcal{D}\mathsf{isk}_{n/U_{I}}^{\leq k}}F\left(V\right)\to\operatorname{colim}_{V\in\mathcal{D}\mathsf{isk}_{n/U}^{\leq k}}F\left(V\right)$
induced by the functor $\phi:\int^{I\in\mathcal{I}}\mathcal{D}\mathsf{isk}_{n/U_{I}}^{\leq k}\to\mathcal{D}\mathsf{isk}_{n/U}^{\leq k}$. It will therefore suffice to show that $\phi$ is final. For this, we observe that the map $\psi:\int^{I\in\mathcal{I}}\operatorname{Disj}^{\leq k}\left(U_{I}\right)\to\int^{I\in\mathcal{I}}\mathcal{D}\mathsf{isk}_{n/U_{I}}^{\leq k}$ is final (Proposition B.7). Since final maps have the right cancellation property [Cis19, Corollary 4.1.9], it suffices to show that the composite $\phi\circ\psi$ is final. We prove this by using Lemma 1.11: We must show that, for each finite set $S\subset U$, the category $\int^{I\in\mathcal{I}_{S}}\operatorname{Disj}^{\leq k}\left(U_{I}\right)_{S}$ is final. For this, observe that the projection $\int^{I\in\mathcal{I}_{S}}\operatorname{Disj}^{\leq k}\left(U_{I}\right)_{S}\to\mathcal{I}_{S}$ is a cocartesian fibration whose fibers are weakly contractible (as they are cofiltered), and that its base is also weakly contractible by hypothesis. It follows from Quillen’s theorem B (Proposition A.3) that $\int^{I\in\mathcal{I}_{S}}\operatorname{Disj}^{\leq k}\left(U_{I}\right)_{S}$ is also weakly contractible, as claimed.
[…]
@@
[…]
@@
Let $U\subset M$ be an open set homeomorphic to an open subset of $\mathbb{R}^{n}$, where $n=\dim M$. We claim that $U$ is good. Choose an embedding $U\hookrightarrow\mathbb{R}^{n}$ and identify $U$ with a subset of $\mathbb{R}^{n}$. Let $\operatorname{Cube}\left(U\right)$ denote the poset of open subsets of $U$ that are finite disjoint unions of open cubes lying in $U$, and let $\overline{\operatorname{Cube}\left(U\right)}$ denote the poset of open subsets of elements of $\operatorname{Cube}\left(U\right)$. Since $\operatorname{Cube}\left(U\right)$ is closed under finite intersection, the inclusion $\operatorname{Cube}\left(U\right)\to\overline{\operatorname{Cube}\left(U\right)}$ is final. Moreover, $\overline{\operatorname{Cube}\left(U\right)}$ is a covering sieve of $U$ in the Weiss topology on $\operatorname{Open}\left(M\right)$. Since every element of $\operatorname{Cube}\left(U\right)$ is good, it follows that $\alpha_{U}$ is an equivalence.
@@
Remark 1.16.
Using Variant A.10, we can show that everything in this section is valid in the smooth case if we make the following replacements:
  • manifolds by smooth manifolds;
  • isotopy equivalences by smooth isotopy equivalences;
A similar remark applies to the PL case.
Remark 1.17.
Let $M$ be a manifold, and let $\mathcal{C}$ be a locally small $\infty$-category with small colimits. Isotopy invariance, exhaustivity, and $k$-excisivity of a functor $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ can be detected jointly by the hom-functors (i.e., by the composites $\operatorname{Open}\left(M\right)^{\mathrm{op}}\xrightarrow{F}\mathcal{C}^{\mathrm{op}}\xrightarrow{\mathcal{C}\left(-,C\right)}\mathcal{S}$, where $C$ ranges over the objects in $\mathcal{C}$). Therefore, some of the key results of this section (like Lemma 1.4) could have been proved by outsourcing everything to the case of space-valued presheaves, covered in [Wei99]. We have opted for the current approach because we believe that it is more conceptual, and also because it clarifies the connection between polynomials and Weiss sheaves.
2. Taylor Towers and Their Convergence
In the previous section, we saw that polynomial approximations exist, and that they are constructed by Kan extensions. In this section, we use it to construct the Taylor (co)tower of an isotopy functor and discuss when the tower converges to the original functor.
Let $M$ be a manifold, and let $\mathcal{C}$ be an $\infty$-category with small colimits. Theorem 1.3 says that for each isotopy functor $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ and each $k\geq 0$, there is a best approximation of $F$ by $k$-excisive, exhaustive isotopy functor, namely, the left Kan extension $T_{k}F$ of $F|\operatorname{Disj}^{\leq k}\left(M\right)$. This is called the $k$th polynomial approximation of $F$. We now organize these approximations into a cotower
@@ 2. Taylor Towers and Their Convergence
Let $\mathcal{A}$ be an $\infty$-category equipped with a sequence $\mathcal{A}^{0}\subset\mathcal{A}^{1}\subset\cdots$ of full subcategories. Set $\mathcal{A}^{\infty}=\bigcup_{i}\mathcal{A}^{i}$. We let $\mathbb{T}\left(\mathcal{A}\right)\subset\left(\mathbb{Z}_{\geq 0}\star\{\infty\}\right)\times\mathcal{A}$ denote the full subcategory spanned by the objects $\left(i,A\right)$, where $i\in\mathbb{Z}_{\geq 0}$ and $A\in\mathcal{A}_{i}$, together with the objects $\left(\infty,A\right)$, where $A\in\mathcal{A}$.
Let $\mathcal{C}$ be another $\infty$-category with small colimits, and let $F:\mathcal{A}\to\mathcal{C}$ be a functor. A functor $\mathbb{T}\left(F\right):\left(\mathbb{Z}_{\geq 0}\star\{\infty\}\right)\times\mathcal{A}\to\mathcal{C}$ is called a Taylor cotower of $F$ (with respect to the subcategories $\{\mathcal{A}^{i}\}_{i\geq 0}$) if it is a left Kan extension of the composite
$\mathbb{T}\left(\mathcal{A}\right)\xrightarrow{\text{projection}}\mathcal{A}\xrightarrow{F}\mathcal{C}.$
For each $k\in\mathbb{Z}_{\geq 0}\star\{\infty\}$, we set $T_{k}F=\mathbb{T}\left(F\right)|\{k\}\times\mathcal{A}$. Note that for each $k\in\mathbb{Z}_{\geq 0}$ and $A\in\mathcal{A}$, the inclusion
@@
A Taylor cotower $\mathbb{T}\left(F\right)$ of $F$ is said to be convergent if the functor $\mathbb{Z}_{\geq 0}\star\{\infty\}\to\operatorname{Fun}\left(\mathcal{A},\mathcal{C}\right)$ adjoint to $\mathbb{T}\left(F\right)$ is a colimit diagram.
@@
Let $M$ be a manifold, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ be a functor. A Taylor cotower of $F$ is a Taylor cotower of $F$ with respect to the subcategories $\{\operatorname{Disj}^{\leq i}\left(M\right)\}_{i\geq 0}$. If $F$ is an isotopy functor, then for each $k\in\mathbb{Z}$, the $k$th level $T_{k}F$ of a Taylor cotower of $F$ is polynomial of degree $\leq k$ (Theorem 1.3).
Warning 2.3.
Despite the name, the constituents of Taylor cotowers of non-isotopy functors may fail to be polynomial. For example, consider the functor $F:\operatorname{Open}\left(\mathbb{R}\right)\to\mathsf{Set}$ defined by $F\left(U\right)=U\times U$. It is easy to check that $T_{1}F$ is not excisive.
@@
Let $\mathcal{A}$ be an $\infty$-category, and let $\mathcal{A}_{0}\subset\mathcal{A}_{1}\subset\cdots$ be a sequence of full subcategories of $\mathcal{A}$. Set $\mathcal{A}_{\infty}=\bigcup_{i\geq 0}\mathcal{A}_{i}$. Let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:\mathcal{A}\to\mathcal{C}$ be a functor. The following conditions are equivalent:
@@
We will use the transitivity of Kan extensions [Lur09, Proposition 4.3.2.8]. Let $T=\mathbb{T}\left(F\right)$ be a Taylor cotower of $F$. Since colimits in functor categories can be formed pointwise [Lur25, Tag 02X9], an easy finality argument shows that condition (1) is equivalent to the following condition:
@@
    $T$ is a left Kan extension of $T|\mathbb{Z}_{\geq 0}\times\mathcal{A}$.
Let $\mathcal{X}\subset\left(\mathbb{Z}_{\geq 0}\star\{\infty\}\right)\times\mathcal{A}$ denote the full subcategory spanned by the objects in $\mathbb{Z}_{\geq 0}\times\mathcal{A}$ and the objects in $\{\infty\}\times\mathcal{A}_{\infty}$. Since every object $A\in\mathcal{A}_{\infty}$ belongs to some $\mathcal{A}_{i}$, the functor $T|\mathcal{X}$ is a left Kan extension of $T|\mathbb{Z}_{\geq 0}\times\mathcal{A}$. Therefore, condition (1) is equivalent to the following condition:
@@
On the other hand, $T|\mathcal{X}$ is a left Kan extension of $T|\left(\mathbb{Z}_{\geq 0}\star\{\infty\}\right)\times\mathcal{A}_{\infty}$, so (1′′) is equivalent to the following condition:
@@
    $T$ is a left Kan extension of $T|\left(\mathbb{Z}_{\geq 0}\star\{\infty\}\right)\times\mathcal{A}_{\infty}$.
Since $T|\mathbb{Z}_{\geq 0}\times\mathcal{A}$ is a left Kan extension of $T|\mathbb{Z}_{\geq 0}\times\mathcal{A}_{\infty}$, condition (1′′′) is equivalent to condition (2), and we are done. ∎
@@
Let $M$ be a manifold, and let $\mathcal{C}$ be an $\infty$-category with small colimits. An isotopy functor $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ has a convergent Taylor cotower if and only if $F$ satisfies the equivalent conditions of Theorem 1.10.
@@
When we can talk about connectivity of morphisms, we can give a better convergence result. We will focus on one such instance, where the target $\infty$-category is a stable $\infty$-category with t-structures. (See Section B.3 for a brief review of t-structures.) The following definition follows [GW99, Theorem 2.3].
Definition 2.6.
Let $M$ be a smooth manifold, let $\mathcal{C}$ be a stable $\infty$-category with $t$-structures, let $F:\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}$ be an exhaustive isotopy functor, and let $\rho$ and $c$ be integers. We say that $G$ is $\rho$-analytic with excess $c$ if it satisfies the following condition:
  • ($\ast$)
    Let $P\subset M$ be a smooth compact submanifold with boundary, and let $Q_{0},\dots,Q_{r}\subset M\setminus\operatorname{Int}\left(P\right)$ be pairwise disjoint, compact, codimension 0 smooth submanifolds with boundary. Suppose that each $Q_{i}$ has handle index $q_{i}<\rho$ (i.e., it is obtained from $\partial Q_{i}$ by attaching handles of indices $\leq q_{i}$). Then the $\left(r+1\right)$-cube
    $\mathcal{P}\left(\{0,\dots,r\}\right)^{\mathrm{op}}\to\mathcal{C},\,S\mapsto F\left(\operatorname{Int}\left(P\cup\bigcup_{s\in S}Q_{s}\right)\right)$
    is $\left(c+\sum_{i=0}^{r}\left(\rho-q_{i}\right)\right)$-cartesian.
The following result should be compared with [GW99, Theorem 2.3]. It also generalizes [RW14, Theorem E.5].
Theorem 2.7.
Let $\mathcal{C}$ be a stable $\infty$-category with a t-structure, let $M$ be a smooth manifold, let $\rho$ and $c$ be integers, and let $F:\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}$ be an exhaustive isotopy functor. Suppose that the following conditions are satisfied:
  1. (1)
    $F$ is $\rho$-analytic with excess $c$.
  2. (2)
    $\mathcal{C}$ has small limits.
  3. (3)
    The functor $\pi_{0}:\mathcal{C}\to\mathcal{C}^{\heartsuit}$ preserves countable products.
Then for every $k\geq 1$ and every open set $W\subset M$ admitting a proper Morse function whose critical points have indices less than $q$, the map
$\eta_{k-1,W}:F\left(W\right)\to T_{k-1}F\left(W\right)$
is $\left(c+k\left(\rho-q\right)-1\right)$-connected. [^2] This is slightly weaker than the estimate of [GW99], where it says $c+k\left(\rho-q\right)$ instead of $c+k\left(\rho-q\right)-1$. However, a close inspection of the proof of loc. cit. shows that it only gives the same estimate as ours. Michael Weiss informed me that this issue can probably be fixed, but the details are not written down yet. We believe that his technique will be applicable to our setting too, so we should be able to say that $\eta_{k-1,W}$ is $c+k\left(\rho-q\right)$-connected. In particular, if $\rho>q$ and $\bigcap_{n}\mathcal{C}_{\geq n}=0$, then the map
$F\left(W\right)\to\operatorname{lim}_{k}T_{k}F\left(W\right)$
is an equivalence.
Proof.
We follow [GW99, Theorem 2.3]. Shifting degrees if necessary, we may assume that $c=0$. Suppose first that $\overline{W}$ is a compact smooth submanifold with boundary of $M$ and admits a handle decomposition with handles of indices at most $q$. We will show that $\eta_{k-1,W}$ is $k\left(\rho-q\right)$-connected by induction on $q$ and the number of $q$-handles.
If $q=0$, then $W$ is the disjoint union of open balls, say $W_{1},\dots,W_{l}$. If $l\leq k-1$, then $\eta_{k-1,W}$ is an equivalence, and we are done. If $l\geq k$, it suffices to show that for each $k\leq t\leq l$, the map
$\theta_{t}:T_{t}F\left(W\right)\to T_{t-1}F\left(W\right)$
is $t\rho$-connected. For this, for each subset $S\subset\pi_{0}\left(W\right)$, let $W_{S}\subset W$ denote the union of the elements in $S$. Consider the following commutative diagram
$\coprod_{S\subset\pi_{0}(W)}\mathcal{P}_{0}(S)$$\coprod_{S\subset\pi_{0}(W),\,|S|=t}\operatorname{Disj}^{\leq t-1}(W_{S})$$\operatorname{Disj}^{\leq t-1}(W)$$\coprod_{S\subset\pi_{0}(W)}\mathcal{P}(S)$$\coprod_{S\subset\pi_{0}(W),\,|S|=t}\operatorname{Disj}^{\leq t}(W_{S})$$\operatorname{Disj}^{\leq t}(W),$$\scriptstyle{\phi_{0}}$$\scriptstyle{\phi}$
where the map $\phi|\mathcal{P}\left(S\right)$ is given by $R\mapsto W_{R}$. The right-hand square is a pushout in $\mathcal{C}\mathsf{at}_{\infty}$ (because the Joyal model structure is left proper), and the maps $\phi_{0}$ and $\phi$ are final by inspection. It follows from Remark B.10 that $\theta_{t}$ is a pullback of the map
$\prod_{S\subset\pi_{0}\left(W\right)}G\left(W_{S}\right)\to\prod_{S\subset\pi_{0}\left(W\right)}\operatorname{lim}_{R\in\mathcal{P}_{0}\left(S\right)}G\left(W_{R}\right),$
which is $t\rho$-connected by the analyticity assumption (applied to $P=\emptyset$ and $Q_{i}=\overline{W_{i}}$). Hence $\theta_{t}$ is $t\rho$-connected, as required.
For the inductive step, suppose we can write $\overline{W}=N\cup H$, where $H\cong D^{q}\times D^{n-q}$ is a $q$-handle and $N$ has one less $q$-handle than $\overline{W}$. Choose disjoint closed disks $C_{0},\dots,C_{k}\subset\operatorname{Int}D^{q}$, and set $A_{i}=C_{i}\times D^{n-q}$. By the induction hypothesis, for each $S\in\mathcal{P}_{0}\left(\{0,\dots,k\}\right)$, the map
$G\left(W\setminus\bigcup_{i\in S}A_{i}\right)\to T_{k-1}\left(W\setminus\bigcup_{i\in S}A_{i}\right)$
is $k\left(\rho-q+1\right)$-connected. It follows from Corollary B.15 that the map
$\operatorname{lim}_{S\in\mathcal{P}_{0}\left(S\right)}G\left(W\setminus\bigcup_{i\in S}A_{i}\right)\to\operatorname{lim}_{S\in\mathcal{P}_{0}\left(S\right)}T_{k-1}\left(W\setminus\bigcup_{i\in S}A_{i}\right)\simeq T_{k-1}\left(W\right)$
is $k\left(\rho-q\right)$-connected. Moreover, the cube $S\mapsto G\left(W\setminus\bigcup_{i\in S}A_{i}\right)$ is $k\left(\rho-q\right)$-cartesian by the analyticity assumption. Hence the map $G\left(W\right)\to T_{k-1}\left(W\right)$ is $k\left(\rho-q\right)$-connected (Proposition B.12), completing the induction.
The general case and the last claim are taken care of by the Milnor exact sequence (Lemma B.16; note that conditions (2) and (3) ensure that the hypotheses of the lemma are satisfied), and we are done. ∎